# SurgPointTransformer: Vertebrae Shape Completion with RGB-D Data

### v1

@@ Abstract
State-of-the-art computer- and robot-assisted surgery systems heavily depend on intraoperative imaging technologies such as CT and fluoroscopy to generate detailed 3D visualization of the patient’s anatomy. While imaging techniques are highly accurate, they are based on ionizing radiation and expose patients and clinicians. This study introduces an alternative, utterly radiation-free approach for reconstructing the 3D spine anatomy using RGB-D data. Drawing inspiration from the 3D ”mental map” that surgeons form during surgeries, we introduce SurgPointTransformer, a shape completion approach for surgical applications that can accurately reconstruct the unexposed spine regions from sparse observations of the exposed surface.
Our method involves two main steps: segmentation and shape completion. The segmentation step includes spinal column localization and spinal column segmentation, followed by vertebra-wise segmentation. The segmented vertebra point clouds are then subjected to SurgPointTransformer, which leverages an attention mechanism to learn patterns between visible surface features and the underlying anatomy. For evaluation, we utilize an ex-vivo dataset of nine specimens. Their CT data is used to establish ground truth data that were used to compare to the outputs of our methods. Our method significantly outperforms the state-of-the-art baselines, achieving an average Chamfer Distance of 5.39, an F-Score of 0.85, an Earth Mover’s Distance of 0.011, and a Signal-to-Noise Ratio of 22.90 dB.
This study demonstrates the potential of our reconstruction method for 3D vertebral shape completion. It enables 3D reconstruction of the entire lumbar spine and surgical guidance without ionizing radiation or invasive imaging. Our work contributes to computer-aided and robot-assisted surgery, advancing the perception and intelligence of these systems.
@@ 1 Introduction
Accurate execution in spine surgery is essential due to the spine’s complex anatomy and the proximity of critical structures like the spinal cord, nerves, and major blood vessels. Precision is crucial for minimizing complications and improving patient outcomes. Traditionally, surgeons have used the freehand technique, relying on their deep understanding of anatomical landmarks to avoid damaging vital structures. However, advancements in intraoperative navigation technologies, such as computer-assisted surgery (CAS) and surgical robotics, now assist in key stages of surgery [1]. These technologies help localize the area of interest, verify fracture patterns, and ensure the proper placement of instruments and implants [2]. Fundamental components of these systems are intraoperative imaging technologies, allowing surgeons to monitor anatomical structures and surgical tools in real-time, visualizing unexposed or partially exposed areas like the pedicle region of a vertebra. As a result, procedures are more accurately aligned with preoperative plans.
@@
Only a few studies explored the effectiveness of shape completion approaches in the medical domain. Li et al. [13] developed the Anatomy Completor framework, using a denoising autoencoder (DAE) to reconstruct anatomical shapes from incomplete CT images. Beetz et al. [14] used the Point Completion Network (PCN) to reconstruct cardiac anatomy from 2D MRI slices. Similarly, Gafencu et al. [15] investigated shape completion for spinal anatomy using VRCNet from 3D US reconstructions. To the authors’ knowledge, shape completion performance on RGB-D data of human anatomy has not yet been explored.
Our work addresses this knowledge gap, providing new insights and advancements in shape completion of the human anatomy. This work is inspired by the ability of experienced surgeons to mentally reconstruct unseen anatomical structures by complementing visible anatomy with their knowledge of human anatomy. We hypothesize that the attention mechanism of transformer networks can emulate this behavior by learning patterns and correlations between visible and hidden anatomical structures. Moreover, despite its significant dependence on upstream tasks like segmentation, most state-of-the-art methods treat shape completion as an isolated problem. We propose a fully-fletched pipeline for 3D reconstruction of the complete shape of the spine anatomy directly from raw RGB-D camera data of surgical procedures. Our method involves spinal column localization, spinal column segmentation, vertebra-level segmentation, and shape completion. We evaluated our method on the SpineDepth dataset [16]. Using it as a baseline, we benchmark our method against Gafencu et al.’s [15] work. Our approach eliminates the need for traditional registration processes, eliminates intraoperative radiation exposure, and addresses key concerns in current surgical practices.
@@ 2 Methods
Our pipeline, built around shape completion, comprises spinal column localization, spinal column segmentation, and vertebra-level segmentation as preprocessing steps (see Figure 1). First, we capture color ($I_{RGB}$) and depth data ($I_{Depth}$), as described in Section 2.1. Our method localizes the spinal column with a bounding box ($B_{Spine}$) and segments the spinal column by generating a spine segmentation mask ($M_{Spine}$) and then its point cloud ($PCD_{Spine}$). It is followed by a vertebra-wise segmentation resulting in color-coded point cloud ($Pred_{Seg}$), detailed in Section 2.2. These vertebra segments ($Pred_{Partial}$) are then input to our shape completion network SurgPointTransformer (Section 2.3). The predicted point clouds ($Pred_{Complete}$) are then converted to 3D meshes ($Pred_{3D}$) using Poisson surface reconstruction. We detail the evaluation metrics for assessing the method’s effectiveness in Section 2.4. All training and evaluations were performed using an NVIDIA Tesla V100 GPU with 16GB RAM and Python 3.11.5 and PyTorch 2.4.0.
Figure 1: Pipeline of the proposed method for vertebrae shape completion from RGB-D data. A - Segmentation Module: The color image ($I_{RGB}$) from the RGB-D data is fed into the spine detector, which localizes the spine’s position with a bounding box $B_{Spine}$ (indicated in fuchsia). $B_{Spine}$ and $I_{RGB}$ are then passed to the spine segmentation model, resulting in spine point cloud ($PCD_{Spine}$). The vertebra-level segmentation module produces color-coded segmentation ($Pred_{Seg}$) for each vertebra level, where red, green, blue, yellow, fuchsia, and black colors correspond to L1, L2, L3, L4, L5, and background, respectively. B - Shape Completion Module: The segmented vertebra point clouds ($Pred_{Partial}$) are input to our SurgPointTransformer, which reconstructs the complete shape of each vertebra ($Pred_{Complete}$). C - Poisson Surface Reconstruction: The completed point clouds are converted into 3D meshes ($Pred_{3D}$).
@@ 2.2 Segmentation
Spinal Column Localization: We trained a YOLOv8 network [17] to detect the spine region in color images. YOLOv8 outputs a bounding box ($B_{Spine}$) around the detected spine region. The ground truth data for training were generated automatically from 3D vertebra meshes. We applied a respective transformation matrix ($T_{CT}^{CAM}$), available through the calibration chain described in [18], to convert these meshes from CT to camera space to align them with the RGB-D data. Using the camera’s intrinsic parameters, we rendered a binary mask of the lumbar spine from the 3D meshes. The bounding box of this binary mask was used as ground truth for training YOLOv8. The model was trained separately for each fold, with training limited to three epochs.
@@
Each point in $PCD_{Spine}$ was labeled with different vertebral levels, including L1 to L5 and background, to train the multi-class point cloud segmentation network. This process was similar to the ground truth generation for the spinal column localization. Using aligned 3D meshes with the RGB-D data and the camera’s intrinsic parameters, we rendered each vertebra level to label each pixel on $I_{RGB}$. We applied the resulting vertebra level label mask ($M_{Vert}$) on $PCD_{Spine}$ to generate ground truth segmentation labels ($GT_{Seg}$). We downsampled $PCD_{Spine}$ to 10’000 points using open3d 0.18.0 [22] library to reduce computational complexity during training. However, the full-resolution, dense point cloud was used during inference.
@@ 2.4 Evaluation
We conducted a statistical analysis to compare the performance of segmentation methods (using XYZ vs. XYZRGB input configurations) and shape completion methods (SurgPointTransformer vs. VRCNet). Kolmogorov-Smirnov test showed that the data did not follow a normal distribution. As a result, non-parametric testing with the Mann-Whitney U test was applied. A p-value of less than 0.05 was considered statistically significant. We also examined the correlations between the data characteristics (specimen, vertebra level) and localization and segmentation performance on the overall effectiveness of our pipeline, using the Pearson correlation method to assess the strength and direction of the relationships.
@@ 3 Results
Table 1: Shape Completion Results for our approach against VRCNet. This table shows the performance metrics for shape completion, including IoU, CD, F-Score, EMD, CD_top, CD_bottom, and SNR. Results are provided for each vertebra level and their averages. Statistically significant higher values between the two methods are highlighted in bold. $\mu$ means the average value over all vertebrae levels.
@@
Table 2: The vertebra-wise segmentation results: IoU and Accuracy for each class and overall, comparing two different input configurations. Statistically significant higher values comparing two input configurations are highlighted in bold.
@@
Figure 4: Segmentation results for XYZ and XYZRGB inputs are shown with ground truth (GT). Point clouds are overlaid on 3D vertebra meshes, with L1 (red), L2 (green), L3 (blue), L4 (yellow), and L5 (pink) highlighted separately.
Table 3: Performance Metrics for Segmentation and Shape Completion Tasks Using Our Proposed Method. This table displays the results for each specimen, including IoU_input, IoU_seg, and Accuracy for the segmentation task, as well as CD, F-Score, EMD, CD_top, CD_bottom, and SNR for the shape completion task. Statistically significant higher values for each specimen are highlighted in black bold, while the least are highlighted in blue bold.

### v2

@@ Abstract
State-of-the-art computer- and robot-assisted surgery systems heavily depend on intraoperative imaging technologies such as CT and fluoroscopy to generate detailed 3D visualization of the patient’s anatomy. While imaging techniques are highly accurate, they are based on ionizing radiation and expose patients and clinicians. This study introduces an alternative, radiation-free approach for reconstructing the 3D spine anatomy using RGB-D data. Drawing inspiration from the 3D ”mental map” that surgeons form during surgeries, we introduce SurgPointTransformer, a shape completion approach for surgical applications that can accurately reconstruct the unexposed spine regions from sparse observations of the exposed surface.
Our method involves two main steps: segmentation and shape completion. The segmentation step includes spinal column localization and segmentation, and vertebra-wise segmentation. The segmented vertebra point clouds are then subjected to SurgPointTransformer, which leverages an attention mechanism to learn patterns between visible surface features and the underlying anatomy. For evaluation, we utilize an ex-vivo dataset of nine specimens. Their CT data is used to establish ground truth data that were used to compare to the outputs of our methods. Our method significantly outperforms the state-of-the-art baselines, achieving an average Chamfer Distance of 5.39, an F-Score of 0.85, an Earth Mover’s Distance of 0.011, and a Signal-to-Noise Ratio of 22.90 dB.
This study demonstrates the potential of our reconstruction method for 3D vertebral shape completion without ionizing radiation or invasive imaging. Our work contributes to computer-aided and robot-assisted surgery, advancing the perception and intelligence of these systems.
@@ 1 Introduction
Accurate surgical execution in spine surgery is essential due to the spine’s complex anatomy and the proximity of critical structures like the spinal cord, nerves, and major blood vessels. Precision is crucial for minimizing complications and improving patient outcomes. Traditionally, surgeries are performed free-hand, where surgeons rely on their deep understanding of anatomy to avoid damaging vital structures. However, advancements in intraoperative navigation technologies, such as computer-assisted surgery (CAS) and surgical robotics, have been gradually integrated into the surgical pipeline [1]. These technologies help localize the area of interest, verify fracture patterns, and ensure and the proper execution of the preoperative plan [2]. Fundamental components of these systems are intraoperative imaging technologies, allowing surgeons to monitor anatomical structures and surgical tools, visualizing unexposed or partially exposed areas like the pedicle region of a vertebra. As a result, procedures are more accurately aligned with preoperative plans.
@@
Only a few studies explored the effectiveness of shape completion approaches in the medical domain. Li et al. [13] developed the Anatomy Completor framework, using a denoising autoencoder (DAE) to reconstruct anatomical shapes from incomplete CT images. Beetz et al. [14] used the Point Completion Network (PCN) to reconstruct cardiac anatomy from 2D MRI slices. Similarly, Gafencu et al. [15] investigated shape completion for spinal anatomy using VRCNet from 3D US reconstructions. To the best of our knowledge, shape completion performance on RGB-D data of human anatomy has not yet been explored.
Our work addresses this knowledge gap, providing new insights and advancements in shape completion of the human anatomy. This work is inspired by the ability of experienced surgeons to mentally reconstruct unseen anatomical structures by complementing visible anatomy with their knowledge of human anatomy. We hypothesize that the attention mechanism of transformer networks can emulate this behavior by learning patterns and correlations between visible and hidden anatomical structures. Moreover, despite its significant dependence on upstream tasks like segmentation, most state-of-the-art methods treat shape completion as an isolated problem. We propose a fully-fletched pipeline for 3D reconstruction of the complete shape of the spine anatomy directly from raw RGB-D camera data of surgical procedures. Our method involves spinal column localization, spinal column segmentation, vertebra-level segmentation, and shape completion. We evaluated our method on the SpineDepth dataset [16] and using it as a baseline, and benchmarked it against Gafencu et al.’s [15] work. Our approach eliminates the need for traditional registration processes, intraoperative radiation exposure, and addresses key concerns in current surgical practices.
@@ 2 Methods
Our pipeline, built around shape completion, comprises spinal column localization, spinal column segmentation, and vertebra-level segmentation as preprocessing steps (see Figure 1). First, we capture color ($I_{RGB}$) and depth data ($I_{Depth}$), as described in Section 2.1. Our method localizes the spinal column with a bounding box ($B_{Spine}$) and segments the spinal column by generating a spine segmentation mask ($M_{Spine}$) and then applies $M_{Spine}$ on the RGB-D data to produce spinal point cloud ($PCD_{Spine}$). It is followed by a vertebra-wise segmentation resulting in color-coded point cloud ($Pred_{Seg}$), detailed in Section 2.2. The resulting vertebra segments ($Pred_{Partial}$) are then input to our shape completion network SurgPointTransformer (Section 2.3). The predicted point clouds ($Pred_{Complete}$) are then converted to 3D meshes ($Pred_{3D}$) using Poisson surface reconstruction. We detail the evaluation metrics for assessing the method’s effectiveness in Section 2.4. All training and evaluations were performed using an NVIDIA Tesla V100 GPU with 16GB RAM and Python 3.11.5 and PyTorch 2.4.0.
Figure 1: Pipeline of the proposed method for vertebrae shape completion from RGB-D data: $I_{RGB}$ from the RGB-D data is fed into the spinal column detector, which localizes the spine’s position with a bounding box $B_{Spine}$ (indicated in fuchsia). $B_{Spine}$ and $I_{RGB}$ are then passed to the spine segmentation model, resulting in a spinal segmentation mask ($M_{Spine}$).This mask then applied to $I_{RGB}$ and $I_{Depth}$ to produce $PCD_{Spine}$. The vertebra-level segmentation module produces color-coded segmentation ($Pred_{Seg}$) for each vertebra level, where red, green, blue, yellow, fuchsia, and black colors correspond to L1, L2, L3, L4, L5, and background, respectively. The segmented vertebra point clouds ($Pred_{Partial}$) are input to our SurgPointTransformer, which reconstructs the complete shape of each vertebra ($Pred_{Complete}$). The completed point clouds are converted into 3D meshes ($Pred_{3D}$).
@@ 2.2 Segmentation
Spinal Column Localization: We trained a YOLOv8 network [17] to detect the spine region in color images. YOLOv8 outputs a bounding box ($B_{Spine}$) around the detected spine region. The ground truth data for training were generated automatically from 3D vertebra meshes ($Mesh_{Vert}$). We applied a respective transformation matrix ($T_{CT}^{CAM}$), available through the calibration chain described in [18], to convert $Mesh_{Vert}$ from CT to camera space to align them with the RGB-D data. Using the camera’s intrinsic parameters, we rendered a vertebra-level mask ($M_{Vert}$) of the lumbar spine from the 3D meshes. The bounding box of $M_{Vert}$ was used as ground truth for training YOLOv8. The model was trained separately for each fold, with training limited to three epochs.
@@
Each point in $PCD_{Spine}$ was labeled with different vertebral levels, including L1 to L5 and background, to train the multi-class point cloud segmentation network. We applied $M_{Vert}$ on $PCD_{Spine}$ to generate ground truth segmentation labels ($GT_{Seg}$). We downsampled $PCD_{Spine}$ to 10’000 points using open3d 0.18.0 [22] library to reduce computational complexity during training. However, the full-resolution, dense point cloud was used during inference.
@@ 2.4 Evaluation
We also examined the correlations between the data characteristics (specimen, vertebra level) and localization and segmentation performance on the overall effectiveness of our pipeline, using the Pearson correlation method to assess the strength and direction of the relationships.
@@ 3 Results
Table 1: Shape Completion Results for our approach against VRCNet. This table shows the performance metrics for shape completion, including IoU, CD, F-Score, EMD, CD_top, CD_bottom, and SNR. Results are provided for each vertebra level and their averages. The superior results between the two methods are highlighted in bold. $\mu$ means the average value over all vertebrae levels.
@@
Table 2: The vertebra-wise segmentation results: IoU and Accuracy for each class and overall, comparing two different input configurations. The best results comparing two input configurations are highlighted in bold.
@@
Figure 4: Segmentation results for XYZ and XYZRGB inputs are shown with ground truth. Point clouds are overlaid on 3D vertebra meshes, with L1 (red), L2 (green), L3 (blue), L4 (yellow), and L5 (pink) highlighted separately.
Table 3: Performance Metrics for Segmentation and Shape Completion Tasks Using Our Proposed Method. This table displays the results for each specimen, including IoU_input, IoU_seg, and Accuracy for the segmentation task, as well as CD, F-Score, EMD, CD_top, CD_bottom, and SNR for the shape completion task. The superior results for each specimen are highlighted in black bold, while the least are highlighted in blue bold.
@@ 5 Conclusion
6 Acknowledgements
This work was supported by the FAROS project and has received funding from the European Union’s Horizon 2020 research and innovation program under grant agreement No. 101016985.